OpenStudy (anonymous):

how do I simplify (-4p)^2

OpenStudy (anonymous):

\[(-4p)^2 \implies (-4)^2 \times (p)^2\]

mathslover (mathslover):

\[\large{(ab)^2=a^2\times b^2}\] \[\large{(-4p)^2=(-4)^2\times (p)^2 }\]can u solve it now?

OpenStudy (jiteshmeghwal9):